# On the consistency of non-commutative geometry inspired Reissner-Nordström black hole solution

### v1

@@ Abstract
where $D$ is the dimension of the spacetime. The point source is smeared by a Gaussian distribution with a width $\sqrt{\theta}$, where $\theta$ is a constant with dimension length squared. One now has a single parameter $\theta$ characterizing the non-commutative effects because in order to preserve Lorentz invariance and unitarity, one should take the anti-symmetric tensor in (1) as $\theta^{\mu\nu}=\theta\operatorname{diag}\left(\epsilon_{ij},\epsilon_{ij},\ldots\right)$.
@@
To set the stage, let us review the non-commutative geometry inspired Schwarzschild black hole. Taking a line element of the following form
@@
For the noncommutative geometry inspired version, we need to modify the energy-momentum tensor by using the substitution rule. The new energy-momentum tensor is given by
@@
$\epsilon|_{\text{matt.}}=\frac{M}{(4\pi\theta)^{3/2}}$ $e^{-r^{2}/4\theta},\quad p_{r}|_{\text{matt.}}=-\epsilon|_{\text{matt.}},\quad p_{\perp}|_{\text{matt.}}=-\epsilon|_{\text{matt.}}-\frac{r}{2}\epsilon|_{\text{matt.}}^{\prime},$ (13)
@@
Now we can turn our attention to Einstein-Maxwell theory whose field equations are given by
@@
$\epsilon+p_{r}+2p_{\perp}\geq 0\quad\Rightarrow\quad p_{\perp}\geq 0$ (34)
since $p_{r}=-\epsilon$. Therefore, the improved energy-momentum tensor that we propose predicts a different region where the energy condition is violated. In Figure 1, we present our results for $Q=5\sqrt{\theta}$. In Figure 1(a), we show that the mass of the black hole should satisfy $M>M_{0}=1.79\sqrt{\theta}$ in order to have two distinct event horizons $r_{-}$ and $r_{+}$. In Figure 1(b), the strong energy curves for the standard and the improved energy-momentum tensors are shown with the choice $M=2\sqrt{\theta}$, for which the inner horizon is located at $r=r_{-}=2.32\sqrt{\theta}$ and the outer is at $r=r_{+}=4.19\sqrt{\theta}$. For the former case, the energy condition is violated when $r<r_{*}=0.61\sqrt{\theta}$. However, with the improved energy-momentum tensor, the violation occurs when $r<r_{*}=2.95\sqrt{\theta}$. For the violation to be hidden inside a Cauchy horizon, we need to have $r_{*}<r_{-}$. Therefore, according to the standard energy-momentum tensor, which does not satisfy all the components of the field equations, the violation occurs inside the Cauchy horizon. However, there is violation outside the Cauchy horizon according to the improved energy-momentum tensor.
@@
This result suggests that the causal structure of non-commutative geometry inspired charged black holes may need to be reconsidered whenever the point electric charge is smeared. Since our improvement yields violations of the energy conditions outside the Cauchy horizon for some values of parameters, it might have important implications for the validity of cosmic censorship hypothesis and for potential observational signatures. As an alternative way to get rid of the inconsistency of Einstein equations without any manipulation by hand, one can also try to regularize the electric field by using a non-linear electromagnetic coupling. There are quite a number of examples in the literature including the celebrated Born-Infeld electrodynamics that give a regular electric field for a point charge source. Although contradicting the main philosophy of non-commutative geometry inspired gravitational solutions, which is using the substitution (4) for any point sources, this seems to be an interesting possibility. We hope to report on this issue in future work.

### v2

@@ Abstract
where $D$ is the dimension of the spacetime. The point source is smeared by a Gaussian distribution with a width $\sqrt{\theta}$, where $\theta$ is a constant with dimension length squared. One now has a single parameter $\theta$ characterizing the non-commutative effects because in order to preserve Lorentz invariance and unitarity, one should take the anti-symmetric tensor in (1) as $\theta^{\mu\nu}=\theta\operatorname{diag}\left(\epsilon_{ij},\epsilon_{ij},\ldots\right)$ [13].
@@
To set the stage, let us review the non-commutative geometry inspired Schwarzschild black hole [14]. Taking a line element of the following form
@@
For the noncommutative geometry inspired version, we need to modify the energy-momentum tensor by using the substitution rule (4). The new energy-momentum tensor is given by
@@
$\epsilon|_{\text{matt.}}=\frac{M}{(4\pi\theta)^{3/2}}$ $e^{-r^{2}/4\theta},\quad p_{r}|_{\text{matt.}}=-\epsilon|_{\text{matt.}},\quad p_{\perp}|_{\text{matt.}}=-\epsilon|_{\text{matt.}}-\frac{r}{2}\epsilon|_{\text{matt.}}^{\prime}.$ (13)
@@
We can now turn our attention to Einstein-Maxwell theory whose field equations are given by
@@
$\epsilon+p_{r}+2p_{\perp}\geq 0\quad\Rightarrow\quad p_{\perp}\geq 0,$ (34)
since $p_{r}=-\epsilon$. Therefore, the improved energy-momentum tensor that we propose predicts a different region where the energy condition is violated. In Figure 1, we present our results for $Q=5\sqrt{\theta}$. In Figure 1(a), we show that the mass of the black hole should satisfy $M>M_{0}=1.79\sqrt{\theta}$ in order to have two distinct event horizons $r_{-}$ and $r_{+}$. In Figure 1(b), the strong energy curves for the standard and the improved energy-momentum tensors are shown with the choice $M=2\sqrt{\theta}$, for which the inner horizon is located at $r_{-}=2.32\sqrt{\theta}$ and the outer is at $r_{+}=4.19\sqrt{\theta}$. For the former case, the energy condition is violated when $r<r_{*}=0.61\sqrt{\theta}$. However, with the improved energy-momentum tensor, the violation occurs when $r<r_{*}=2.95\sqrt{\theta}$. For the violation to be hidden inside a Cauchy horizon, we need to have $r_{*}<r_{-}$. Therefore, according to the standard energy-momentum tensor, which does not satisfy all the components of the field equations, the violation occurs inside the Cauchy horizon. However, there is violation outside the Cauchy horizon according to the improved energy-momentum tensor. This clearly shows that certain parameter choices, although seemingly safe under the standard treatment, should be avoided to ensure the global hyperbolicity of the solution.
@@
This result suggests that the causal structure of non-commutative geometry inspired charged black holes may need to be reconsidered whenever the point electric charge is smeared. Since our improvement yields violations of the energy conditions outside the Cauchy horizon for some values of parameters, it might have important implications for potential observational signatures. As an alternative way to get rid of the inconsistency of Einstein equations without any manipulation by hand, one can also try to regularize the electric field by using a non-linear electromagnetic coupling. There are quite a number of examples in the literature including the celebrated Born-Infeld electrodynamics that give a regular electric field for a point charge source. Although contradicting the main philosophy of non-commutative geometry inspired gravitational solutions, which is using the substitution (4) for any point sources, this seems to be an interesting possibility. We hope to report on this issue in future work.